cerabean:

Which of the following points lies on the graph of r = 2 - 3cosθ?

NarutoandSasuke:

The choices?

cerabean:

(-1, 0°) (0, 0°) (1, 0°) (2, 0°)

Vocaloid:

for each of these points, the first coordinate is r and the second coordinate is θ simply plug in the given r and θ values to see if the equation is true

Vocaloid:

for example, take the first point (-1, 0°) and let r = -1 and θ be 0 and see if r = 2 - 3cosθ is a true statement

cerabean:

Identify the conic whose equation is given. r=2/3+5 sinθ ellipse parabola hyperbola

NarutoandSasuke:

Post in a new "Ask a question." so you're not warned or suspended..

Vocaloid:

@Zepdrix I don't remember how to do this may I have some assistance?

Zepdrix:

#1\[\large\rm r=\frac23+5\sin \theta\] #2\[\large\rm r=\frac{2}{3+5\sin \theta}\] Is it written like #1 or #2? I just wanna make sure we start from the correct place.

cerabean:

written like#2

Zepdrix:

Let's start by multiplying through by our denominator,\[\large\rm r(3+5\sin \theta)\] Recall that we have this relationship when going between Cartesian and Polar,\[\large\rm r^2=x^2+y^2\]\[\large\rm r=\sqrt{x^2+y^2}\] \[\large\rm x=r \cos \theta\]\[\large\rm y=r \sin \theta\]

Zepdrix:

Woops I made a typo, step one should give us,\[\large\rm r(3+5\sin \theta)=2\]

Zepdrix:

Distributing,\[\large\rm 3r+5r \sin \theta=2\]

Zepdrix:

And then convert to Cartesian,\[\large\rm 3\sqrt{x^2+y^2}+5y=2\]

cerabean:

what would it be though? an ellipse, parabola, or a hyperbola.

Zepdrix:

And do some stuff,\[\large\rm 3\sqrt{x^2+y^2}=(2-5y)\]And square,\[\large\rm 3(x^2+y^2)=(2-5y)^2\] You could go further, but from this point, it looks like we're going to end up with both squared x's and squared y's with different coefficients. So we have an ellipse. There might be an easier way to identify the shape from the start :p but I can't remember...

Zepdrix:

Err wait sorry lemme :)

Zepdrix:

I forgot about the subtraction XD I should be careful haha

cerabean:

I have another question. Write the equation of the conic satisfying the given conditions. focus at the pole, e = 3/4, directrix rsinθ = -2

Zepdrix:

\[\large\rm 3x^2+3y^2 = 4-20y+25y^2\] \[\large\rm 3x^2-22y^2+20y=4\] Ya ya, we're subtracting the y's, completing the square with some value,\[\large\rm 3x^2-(y+c)^2=d\]So we end up with a hyperbola. Ahh math is hard >.<

Zepdrix:

Hmm i dunno bout that next one :d

cerabean:

What is the argument of 3 - 3i? 300° 315° 330° 345°

cerabean:

The modulus of (1 + 2i)^4 is _____. 5 20 25 625

cerabean:

Which of the following is an argument of one of the fourth roots of 16(cos80° + isin80°)?